OpenStudy (anonymous):

Solve 5x +7 = 2x - 3

OpenStudy (anonymous):

minus each side by 2x and seven so that you get.... \[3x=-10\] so your \[x=\frac{-10}{3}\]

OpenStudy (anonymous):

can i get your working please. I feel like youve just given me the answer ..

OpenStudy (anonymous):

\[5x-2x+7-7=2x-2x-3-7\]

OpenStudy (anonymous):

im sorry but im not following you

OpenStudy (anonymous):

other two people viewing this please help me ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

5x+7 = 2x-3 5x = 2x-3-7 ... Subtract 7 from both sides. 5x-2x = -3-7 ... Subtract 2x from both sides. 3x = -10 ... Combine like terms. x = -10/3 ... Divide both sides by 3 to isolate x. So the answer is x = -10/3

OpenStudy (anonymous):

\[5x+7=2x-3\] now add (-7) in both sides \[5x+7-7=2x-3-7\] \[5x=2x-10\] now add -2x in both sides \[5x-2x=2x-10-2x\] \[3x=-10\] now divide both sides by 3 \[\frac{ 3x }{ 3 }= \frac{ -10 }{ 3 }\] \[x= \frac {-10} {3}\]

OpenStudy (anonymous):

so the answer cant be -3.3 ?

OpenStudy (anonymous):

divide -10 in 3, please do it.

OpenStudy (anonymous):

I'm sorry I didn't realise I was on facebook asking for help from retriceoles. I put it in a calculator and thats what I fluttering got. I havent been to school in almost two years I don't fluttering know much so excuse the flutter out of me for being just a little bit stupid. There is no need for you to be a sarcastic flutter !

OpenStudy (anonymous):

really ? I cant even swear. omfg.

OpenStudy (anonymous):

You're trying to combine like terms. It may help you to see it better if you put everything on one side of the equal sign. \[5x+7=2x-3\] Bring the 3 over by adding it to both sides. (-3+3 = 0) \[5x+7+3=2x+0\] Bring the 2x over by subtracting it from both sides (2x-2x=0) \[5x+7+3-2x=0+0\] Now group like terms together. \[(5x-2x)+(7+3)=0\] 5x-2x=3x and 7+3=10 so we get: \[(3x)+(10)=0\] Remember, we are trying to get x by itself, so subtract the 10 from both sides. \[3x=-10\] And finally to get rid of the 3, we divide both sides by 3. (3/3 = 1)\[x=\frac{ -10 }{ 3}\]

OpenStudy (anonymous):

ive got it, thank you ..